OpenStudy (anonymous):

Pythagreon theorem a^2+b^2=c^2 A=x+1 B= x C=X+2. Find all 3 sides in feet

OpenStudy (anonymous):

ok...u just plug the values in the equation and u will get\[(x+1)^2+x^2=(x+2)^2\]\[x^2+2x+1+x^2=x^2+4x+4\]\[x^2-2x-3=0\]this is a quadratic in terms of x...u can factor it or just use direct formula for solving